OpenStudy (anonymous):

write an equation of the line that has slope 2/3 and passes through (1,3)

OpenStudy (anonymous):

3y=2x+7

OpenStudy (aroub):

y=mx+b , m= 2/3 , (1,3) 3=2/3(1)+b 3=2/3 +b 9/3-2/3=b 6/3=b 2=b so the equation is y=2/3x+2